Gilberto and Juanita are farmers. Each one owns a 16-acre plot of land. The following table shows the amount of barley and alfalfa each farmer can produce per year on a given acre. Each farmer chooses whether to devote all acres to producing barley or alfalfa or to produce barley on some of the land and alfalfa on the rest.

Alfalfa (bushels per acre) Barley (bushels per acre)
Gilberto 12 3
Juanita 10 5

1. On the following graph, use the blue line (circle symbol) to plot Gilberto's production possibilities frontier (PPF), and use the purple line (diamond symbol) to plot Juanita's PPF. (insert graph)

2. ___________ has an absolute advantage in the production of alfalfa, and _____________ has an absolute advantage in the production of barley. Gilberto's opportunity cost of producing 1 bushel of barley is ____________ bushels of alfalfa, whereas Juanita's opportunity cost of producing 1 bushel of barley is ________ bushels of alfalfa. Because Gilberto has a ___________ opportunity cost of producing barley than Juanita, ____________ has a comparative advantage in the production of barley, and __________ has a comparative advantage in the production of alfalfa. (fill in the blanks)
Answer #1
Alfalfa (bushels per acre) Barley (bushels per acre)
Gilberto 12*16 = 192 3*16 = 48
Juanita 10*16 = 160 5*16 = 80

Use the information in the above table to plot the graph

Gilberto has an absolute advantage in the production of alfalfa, and Juanita has an absolute advantage in the production of barley. Gilberto's opportunity cost of producing 1 bushel of barley is 12/3 =4 bushels of alfalfa, whereas Juanita's opportunity cost of producing 1 bushel of barley is 10/5 =2 bushels of alfalfa. Because Gilberto has a higher opportunity cost of producing barley than Juanita, Juanita has a comparative advantage in the production of barley, and Gilberto has a comparative advantage in the production of alfalfa